OpenStudy (anonymous):

Given this graph plotting velocity versus time, estimate the acceleration of object A at points X and Y respectively

OpenStudy (irishboy123):

what graph?!

OpenStudy (kayliwarden):

velocity/time should give you your acceleration. \[\left(\begin{matrix} \Delta V \\ \Delta t\end{matrix}\right) = a\]

OpenStudy (anonymous):

Depending on the graph shape, you can estimate the average acceleration by using a secant line between two end points. \[\large \frac{\Delta v}{\Delta t}=a_\text{average}\] Another way to write this is by using the equation of a secant line if you're given a specific function for velocity:\[\large \frac{v(b)-v(a)}{b-a}=a_\text{average}\]

OpenStudy (anonymous):

I should say the `slope` of the secant line*